# Disentangled and Self-Explainable Node Representation Learning

### v1

@@ 4 Experiments
We conduct large-scale experiments to answer the following research questions.
  • Human understandability: How comprehensible and sparse are the explanation substructures generated by DiSeNE?
  • Structural disentanglement: Do the disentangled subgraphs reveal intrinsic properties of node embeddings, like feature correlations and latent positions?
  • Utility for downstream tasks: Are the identified substructures plausible and coherent enough to serve as explanations in downstream tasks?
To address our research questions, we extract topological components of multiple embedding methods, trained on different graph-structured data, by computing edge subsets defined in Eq (2), and analyzing embedding metrics defined in Section 3.3. In the following sections, we describe the data, models, the experimental setup and results from the comparison.
@@ 4.1 Datasets and competitors
We ran experiments on four real-world datasets (Cora, Wiki, FaceBook, PPI), and four synthetic datasets (RingOfCliques, StochasticBlockModel, BA-Cliques and ER-Cliques). RingOfCliques and StochasticBlockModel are implemented in the NetworkX library [^4] https://networkx.org/documentation/stable/reference/generators.html. Statistics for these datasets are provided in Table A1 in the Appendix. Additionally we employ several biological datasets (see Appendix A.3) for the evaluation on multi-label node classification). BA-Cliques is a variation of the BA-Shapes [YBY+19] where we randomly attach cliques instead of house motifs to a Barabási-Albert graph. Similarly, ER-Cliques is generated by attaching cliques to an Erdős-Rényi random graph. All synthetic graphs contain 32 cliques with 10 nodes each as ground-truth sub-structures. For synthetic data, we present only results for plausibility metrics, leaving the other findings in the Appendix A.4.
@@ Methods.
We compare different node embedding methods. Competitors include DeepWalk [PARS14], Graph Autoencoder (GraphAE) [SHV21], InfWalk [CM20], and GraphSAGE [HYL17]. Moreover, we apply the Dine retrofitting approach [PKPA24] to post-process embeddings from DeepWalk and GraphAE. We evaluate our method DiSeNE in two variants: a 1-layer fully-connected encoder (DiSe-FCAE) and a 1-layer convolutional encoder (DiSe-GAE). GNN-based methods are trained using the identity matrix as node features. Details on hyperparameters and training settings are provided in the Appendix A.2.
@@
In experiments on real-world graphs, we investigate latent space interpretability and disentanglement metrics by keeping the output embedding dimension fixed at 128. This dimensionality was chosen to ensure that all methods achieve optimal performance in terms of test accuracy, specifically for link prediction (see the Appendix A.3 for extensive downstream task results). For synthetic data, since we investigated plausibility metric referred to a downstream classifier, thus we did not focus on a specific dimension but we selected the best score metric varying the output dimensions in the list $[2,4,8,16,32,64,128]$. Each reported result is an average over 5 runs. For link prediction, we use a 90%/10% train/test split, and for node classification, we use an 80%/20% split. All results refer to the training set, except for downstream task experiments, where we present results for the test set.
@@ 4.2 Results and Discussion
Are the topological substructures both comprehensible and sparse to support human understandability?
Table 1: Comprehensibility and sparsity results for real-world datasets. Best scores are in bold, second best results are underlined.
Method Comprehensibility Sparsity
@@
DeepWalk .363$\pm$.003 .356$\pm$.002 .602$\pm$.004 .281$\pm$.002 .183$\pm$.001 .165$\pm$.002 .130$\pm$.004 .136$\pm$.003
GraphAE .299$\pm$.001 .248$\pm$.002 .481$\pm$.014 .263$\pm$.003 .182$\pm$.002 .164$\pm$.004 .154$\pm$.003 .135$\pm$.003
InfWalk .281$\pm$.002 .288$\pm$.001 .658$\pm$.006 .312$\pm$.002 .211$\pm$.003 .185$\pm$.002 .318$\pm$.010 .177$\pm$.002
GraphSAGE .358$\pm$.004 .307$\pm$.007 .583$\pm$.003 .306$\pm$.005 .189$\pm$.001 .189$\pm$.003 .145$\pm$.002 .172$\pm$.001
DW+Dine .511$\pm$.051 .496$\pm$.014 .813$\pm$.025 .569$\pm$.022 .317$\pm$.036 .266$\pm$.007 .226$\pm$.009 .188$\pm$.002
GAE+Dine .569$\pm$.004 .591$\pm$.004 .843$\pm$.005 .484$\pm$.007 .290$\pm$.001 .252$\pm$.001 .195$\pm$.002 .198$\pm$.002
DiSe-FCAE .822$\pm$.001 .755$\pm$.003 .971$\pm$.001 .484$\pm$.001 .504$\pm$.001 .419$\pm$.001 .297$\pm$.002 .282$\pm$.001
DiSe-GAE .834$\pm$.003 .762$\pm$.004 .967$\pm$.001 .515$\pm$.001 .496$\pm$.001 .418$\pm$.001 .304$\pm$.003 .254$\pm$.002
Here we explore how well the represented topological structures can serve as global explanations for the node embeddings, quantifying the Comprehensibility in the terms of associations between model parameters and human-understandable units of the input graph, as well as the Sparsity of these associations. In Table 1 we show compact scores as the average values $\frac{1}{K}\sum_{d=1}^{K}\mathrm{Comp}(\mathbf{M}_{d})$ and $1-\frac{1}{K}\sum_{d=1}^{K}\mathrm{Sp}(\mathbf{M}_{d})$ over all the embedding features. For sparsity we report the value subtracted from 1 to have all the scores better with higher values.
DeepWalk and InfWalk show moderate performance in Comprehensibility, excelling slightly on FB but underperforming on PPI, while GraphAE consistently lags behind, particularly on Wiki and PPI. GraphSAGE shows good comprehensibility across Cora and FB. Incorporating DINE improves results, especially for GAE+Dine, which achieves improved scores on all datasets. The proposed models, DiSe-FCAE and DiSe-GAE, deliver the highest overall performance. DiSe-FCAE performs well on FB, while DiSe-GAE excels across Cora and Wiki. However, both models show sub-optimal results on PPI, suggesting potential for further improvement on this dataset.
DeepWalk and GraphAE offer moderate Sparsity, peaking on Cora, but underperform on other datasets. InfWalk excels on FB but shows moderate results elsewhere, while GraphSAGE performs poorly in terms of sparsity across all datasets. DeepWalk and GAE significantly improve their sparsity with Dine, particularly on Cora. For the proposed models, DiSe-FCAE performs best across datasets Cora, Wiki and PPI. Meanwhile, DiSe-GAE obtains the highest value on FB.
Can the identified subgraphs explain the intrinsic characteristics of the node embeddings?
Here we explore how well the defined topological units represent information in the node embedding space, providing insights into how the relative and absolute positioning of topological structures influences the feature encoding within a graph. By quantifying these relationships, we can better understand the underlying patterns and structural information in the graph, potentially leading to more robust and interpretable models. In Table 2 we report Positional Coherence and Overlap Consistency for the examined embedding methods. For the second metric, as node proximity we used the inverse of the shortest path distance with sum as pooling.
Table 2: Overlap consistency and positional coherence results for real-world datasets. Best scores are in bold, second best result are underlined.
@@
DeepWalk .137$\pm$.009 .143$\pm$.006 .115$\pm$.007 .015$\pm$.003 1.078$\pm$0.025 0.835$\pm$0.025 1.119$\pm$0.015 1.009$\pm$0.015
GraphAE .269$\pm$.002 .295$\pm$.004 .273$\pm$.017 .452$\pm$.008 1.023$\pm$0.006 1.040$\pm$0.002 1.001$\pm$0.013 1.016$\pm$0.001
InfWalk .008$\pm$.003 .023$\pm$.002 .021$\pm$.002 .134$\pm$.002 1.004$\pm$0.011 0.998$\pm$0.004 0.938$\pm$0.053 0.999$\pm$0.002
GraphSAGE .211$\pm$.003 .136$\pm$.017 .230$\pm$.007 .097$\pm$.040 1.099$\pm$0.012 1.103$\pm$0.010 1.005$\pm$0.007 1.018$\pm$0.002
DW+Dine .900$\pm$.012 .804$\pm$.032 .851$\pm$.017 .855$\pm$.016 1.790$\pm$0.076 2.126$\pm$0.065 1.792$\pm$0.058 1.247$\pm$0.043
GAE+Dine .560$\pm$.010 .610$\pm$.006 .801$\pm$.016 .646$\pm$.003 2.317$\pm$0.028 2.551$\pm$0.048 1.783$\pm$0.037 1.098$\pm$0.004
DiSe-FCAE .885$\pm$.008 .863$\pm$.006 .939$\pm$.030 .548$\pm$.001 5.210$\pm$0.080 3.540$\pm$0.082 3.348$\pm$0.085 1.283$\pm$0.004
DiSe-GAE .853$\pm$.007 .811$\pm$.008 .889$\pm$.007 .887$\pm$.004 5.300$\pm$0.193 4.343$\pm$0.144 3.388$\pm$0.054 1.261$\pm$0.005
@@
DeepWalk and InfWalk perform poorly for Overlap Consistency, while GraphAE shows moderate scores, particularly on PPI. GraphSAGE performs slightly worse, with the best overlap consistency on FB and Cora. DW+Dine achieves strong scores across all datasets, while GAE+Dine performs solidly but slightly lower, with its best result on FB. The proposed models, DiSe-FCAE and DiSe-GAE, outperform all others, achieving the highest consistency across all datasets except on Cora. DiSe-FCAE excels on FB and Wiki, while DiSe-GAE achieves the best overall score on PPI.
@@
DeepWalk, GraphAE, and GraphSAGE demonstrate moderate Positional Coherence. InfWalk consistently scores around 1.0 on all datasets, indicating stable but unremarkable coherence. Incorporating DINE leads to substantial improvements for both DeepWalk and GAE, achieving notable gains on Cora, Wiki and FB. The proposed models, DiSe-FCAE and DiSe-GAE, far outperform other methods, with DiSe-FCAE achieving top scores on PPI, while DiSe-GAE dominates on Cora, Wiki and FB (though with higher variance): both models show consistent superiority.
@@
Are the identified latent structures sufficiently meaningful to serve as explanations for downstream tasks?
Node embeddings serve as versatile feature representations suitable for downstream tasks, though they typically function as "tabular-like" feature vectors without semantic labels for each feature. This limitation restricts the use of established post-hoc analysis methods [BGG+23] like LIME, SHAP, etc. Our method allows us to link topological substructures with embedding features, thereby assigning semantic labels to node vectors. Consequently, we are able to explain a downstream classifier trained with unsupervised embeddings using feature attribution. Our goal is to assess whether the task-important features align with human understanding by measuring the Plausibility.
In these experiments we consider node classification and link prediction as binary downstream tasks, training a logistic regression classifier $b(x;\boldsymbol{\beta})=\sigma(\sum_{j=1}^{K}\beta_{j}h_{j}(x)+\beta_{0})$, where $x$ is a node/link instance. We use SHAP [LL17] to compute the instance-wise feature attribution values $\{\Psi_{j}(x;b)\}_{j=1\dots K}$. For node classification, we consider positive instances as the nodes inside a clique in the synthetic graph. Accordingly, the ground-truth explanation for a node is the clique it belongs to. For link prediction, we focus on test edges that were inside a clique before removal, where the ground-truth explanation is again the clique itself. We used $f(*)=\mathrm{max(0,*)}$ as non-negative weighting function for computing plausibility scores.
Table 3 shows average Plausibility scores for downstream tasks over test instances. DeepWalk, GraphAE, and InfWalk perform modestly, with DeepWalk scoring the highest among these on Ring-Cl and InfWalk showing relative strength on BA-Cl. GraphSAGE significantly underperforms across all tasks, especially in ER-Cl and BA-Cl. The addition of Dine improves both DeepWalk and GAE. DW+Dine excels with strong performance on Ring-Cl and SBM, while GAE+Dine achieves solid results, particularly on node classification tasks, such as in ER-Cliques. Within the proposed models, DiSe-FCAE and DiSe-GAE, DiSe-FCAE achieves the highest scores on Ring-Cl, BA-Cl, while DiSe-GAE closely follows, leading in SBM and ER-Cl.
@@
Table 3: Plausibility results for synthetic datasets with best scores in bold, second best underlined.
@@
Ring-Cl SBM BA-Cl ER-Cl BA-Cl ER-Cl
DeepWalk .234$\pm$.003 .189$\pm$.003 .173$\pm$.002 .153$\pm$.006 .146$\pm$.002 .141$\pm$.003
GraphAE .183$\pm$.003 .160$\pm$.002 .145$\pm$.004 .145$\pm$.005 .130$\pm$.002 .135$\pm$.006
InfWalk .224$\pm$.005 .181$\pm$.005 .218$\pm$.007 .212$\pm$.008 .129$\pm$.002 .141$\pm$.004
GraphSAGE .252$\pm$.005 .217$\pm$.003 .186$\pm$.006 .178$\pm$.005 .016$\pm$.004 .154$\pm$.002
DW+Dine .943$\pm$.012 .904$\pm$.002 .744$\pm$.008 .722$\pm$.038 .320$\pm$.031 .325$\pm$.007
GAE+Dine .549$\pm$.005 .547$\pm$.014 .418$\pm$.011 .387$\pm$.002 .351$\pm$.011 .394$\pm$.002
DiSe-FCAE .973$\pm$.002 .908$\pm$.006 .943$\pm$.009 .898$\pm$.012 .820$\pm$.011 .791$\pm$.012
DiSe-GAE .969$\pm$.002 .910$\pm$.006 .936$\pm$.003 .941$\pm$.005 .813$\pm$.003 .797$\pm$.009

### v2

@@ 4 Experiments
We conduct extensive experiments to answer the following research questions:
(RQ1) Human understandability: How comprehensible and sparse are the explanation substructures generated by DiSeNE?
(RQ2) Structural disentanglement: Do the disentangled subgraphs reveal intrinsic properties of node embeddings, like feature correlations and latent positions?
(RQ3) Utility for downstream tasks: Are the identified substructures plausible and coherent enough to serve as explanations in downstream tasks?
To address our research questions, we extract topological components from multiple embedding methods, trained on different graph data, by computing edge subsets defined in Eq. (2), and analyzing embedding metrics defined in Section 3.3. In the following, we describe the data, models, experimental setup, and results. Moreover, in the Appendix we report the following supplementary experiments: in section D, we analyze accuracy-interpretability trade-offs for different embedding methods; in section G, we perform ablation studies on the interpretability performances of DiSeNE while varying the entropy regularization and the depth of the convolutional encoder; in section H, we showcase the human-understandability of embeddings in the presence of rich topic and text information in graph data.
@@ 4.1 Datasets and Competitors
We ran experiments on four real-world datasets (Cora, Wiki, FaceBook, PPI), and six synthetic datasets (Ring-of-Cliques, SBM, BA-Cliques, ER-Cliques, Tree-Cliques and Tree-Grids) with planted subgraphs, which serve as ground-truth human rationales for explaining specific node labels. Statistics for these datasets are provided in Table A1 in the Appendix. Additionally we employ several biological datasets (see Appendix C) for the evaluation on multi-label node classification). BA-Cliques and ER-Cliques are variations of the BA-Shapes (Ying et al., 2019) where we randomly attach cliques, instead of house motifs, to Barabási-Albert and Erdős-Rényi random graphs. Tree-Cliques and Tree-Grids (Ying et al., 2019) are composed of a 8-level balanced tree, with cliques and 3x3 grid motifs respectively. Ring-Cliques and SBM (Abbe, 2017) are implemented in NetworkX [^5] https://networkx.org/documentation/stable/reference/generators.html. For synthetic data, we present only results for plausibility metrics, leaving the other findings in the Appendix F.
@@ Methods.
We compare with different node embedding methods in producing explainable feature dimensions. Competitors include shallow encoders DeepWalk (Perozzi et al., 2014), InfWalk (Chanpuriya & Musco, 2020), and deep graph models Graph Autoencoder (GraphAE) (Salha et al., 2020), GraphSAGE (Hamilton et al., 2017) and DGLFRM (Mehta et al., 2019). We also apply the Dine retrofitting approach (Piaggesi et al., 2024) to post-process embeddings from DeepWalk and GraphAE. Moreover, we compare with graph-based local explanation methods GNNExplainer (Ying et al., 2019) and PGExplainer (Luo et al., 2020). Post-hoc explainers are applied to the output of different GNN models for node classification: GCN (Wu et al., 2019), GraphSAGE (Hamilton et al., 2017), and GATv2 (Brody et al., 2022). We evaluate our method DiSeNE in two variants: a 1-layer fully-connected encoder (DiSe-FCAE) and a 1-layer convolutional encoder (DiSe-GAE). GNN-based methods are trained using the identity matrix as node features. Specific training settings are provided in Appendix A.
@@
In experiments on real-world graphs, we investigate latent space interpretability and disentanglement metrics by keeping the output embedding dimension fixed at 128. This dimensionality was chosen to ensure that all methods achieve optimal performance in terms of test accuracy, specifically for link prediction (see the Appendix sections C and D for extensive results on downstream tasks and embedding metrics). For synthetic data, since we investigated plausibility metric referred to a downstream classifier, thus we did not focus on a specific dimension but we selected the best score metric varying the output dimensions in the list $[2,4,8,16,32,64,128]$. Each reported result is an average over 5 runs. For link prediction, we use a 90%/10% train/test split, and for node classification, we use an 80%/20% split. All results refer to the training set, except for downstream task experiments, where we present results for the test set.
@@ 4.2 Results and Discussion
(RQ1) Are the topological substructures both comprehensible and sparse to support human understandability?
Table 1: Topological alignment and sparsity in real data. Best scores in bold, second best underlined.
Method Topological Alignment Sparsity
@@
[…]
Here we explore how well the represented topological structures can serve as global explanations for node embeddings, quantifying the Topological alignment in the terms of associations between model parameters and human-understandable units of the input graph, as well as the Sparsity of these associations. For Topological alignment, we apply modularity maximization to find meaningful clusters (Blondel et al., 2023). In Table 1 we show compact scores as the average values $\frac{1}{K}\sum_{d=1}^{K}\mathrm{Align}(\mathbf{M}_{d})$ and $1-\frac{1}{K}\sum_{d=1}^{K}\mathrm{Sp}(\mathbf{M}_{d})$ over all the embedding features. Since for sparsity we report the value subtracted from 1, all the scores present better results with higher values.
DeepWalk, InfWalk and GraphSAGE show moderate performance in Topological Alignment, excelling slightly on FB but underperforming on PPI, while GraphAE consistently lags behind, particularly on Wiki and PPI. DGLFRM shows good topological alignment across all datasets. Incorporating DINE improves results, especially for GAE+Dine, which achieves improved scores on all datasets. The proposed models, DiSe-FCAE and DiSe-GAE, deliver the highest overall performance. DiSe-FCAE performs well on FB, while DiSe-GAE excels across Cora and Wiki. However, both models show sub-optimal results on PPI, suggesting potential for further improvement on this dataset.
DeepWalk and GraphAE offer moderate Sparsity, peaking on Cora, but underperform on other datasets. InfWalk excels on FB but shows moderate results elsewhere, while GraphSAGE performs poorly in terms of sparsity across all datasets. DeepWalk and GAE significantly improve their sparsity with Dine, particularly on Cora. DGLFRM shows competitive sparsity, excelling in FB and PPI. For the proposed models, DiSe-FCAE and DiSe-GAE perform best across datasets Cora, Wiki.
(RQ2) Can the identified subgraphs explain the intrinsic characteristics of the node embeddings?
Table 2: Overlap consistency and positional coherence in real data. Best scores in bold, second best underlined.
@@
[…]
Here we explore how well the defined topological units represent information in the node embedding space, providing insights into how the relative and absolute positioning of topological structures influences the feature encoding within a graph. By quantifying these relationships, we can better understand the underlying patterns and structural information encoded in graph embeddings. In Table 2 we report Positional Coherence and Overlap Consistency for the examined embedding methods. For the second metric, as node proximity we used the inverse of the shortest path distance with sum as pooling.
@@
DeepWalk, InfWalk and DGLFRM perform poorly for Overlap Consistency, while GraphAE shows moderate scores, particularly on PPI. GraphSAGE performs slightly worse, with the best overlap consistency on FB and Cora. DW+Dine achieves strong scores across all datasets, while GAE+Dine performs solidly but slightly lower, with its best result on FB. The proposed models, DiSe-FCAE and DiSe-GAE, outperform all others, achieving the highest consistency across all datasets except on Cora. DiSe-FCAE excels on FB and Wiki, while DiSe-GAE achieves the best overall score on PPI.
@@
DeepWalk, GraphAE, and GraphSAGE demonstrate moderate Positional Coherence. InfWalk consistently scores around 1.0 on all datasets, indicating stable but unremarkable coherence. Incorporating DINE leads to substantial improvements for both DeepWalk and GAE, achieving notable gains on Cora, Wiki and FB. DGLFRM shows moderate scores, nevertheless scoring the best in PPI. The proposed models, DiSe-FCAE and DiSe-GAE, far outperform other methods, with DiSe-FCAE achieving top scores on PPI, like DGLFRM, while DiSe-GAE dominates on Cora, Wiki and FB (though with higher variance): both models show consistent superiority.
@@
(RQ3) Are the identified latent structures sufficiently meaningful to serve as explanations for downstream tasks?
@@
Table 3: Plausibility for node embeddings in synthetic data. Best scores in bold, second best underlined.
@@
Ring-Cl SBM BA-Cl ER-Cl BA-Cl ER-Cl Tr-Cl Tr-Gr
[…]
Table 4: Plausibility for graph explainers in synthetic data. Best scores in bold, second best underlined.
Node Classif. GnnExplainer PGExplainer DiSeNE
Dataset GCN GSAGE GATv2 GCN GSAGE GATv2 FCAE GAE
[…]
Node embeddings serve as versatile feature representations suitable for downstream tasks, though they typically function as "tabular-like" feature vectors without semantic labels for each feature. This limitation restricts the use of established post-hoc analysis methods (Bodria et al., 2023) like LIME, SHAP, etc. Our method allows us to link topological substructures with embedding features, thereby assigning semantic labels to node vectors. Consequently, we are able to explain a downstream classifier trained with unsupervised embeddings using feature attribution. Our goal is to assess whether the task-important features align with human understanding by measuring the Plausibility.
In these experiments we consider node classification and link prediction as binary downstream tasks, training a logistic regression classifier $b(x;\boldsymbol{\beta})=\sigma(\sum_{j=1}^{K}\beta_{j}h_{j}(x)+\beta_{0})$, where $x$ is a node/link instance. We use SHAP (Lundberg & Lee, 2017) to compute the instance-wise feature attribution values $\{\Psi_{j}(x;b)\}_{j=1\dots K}$. For node classification, we consider positive instances as the nodes inside a clique in the synthetic graph. Accordingly, the ground-truth explanation for a node is the set of nodes within the clique it belongs to. For link prediction, we focus on test edges that were inside a clique before removal, where the ground-truth explanation is again the set of edges inside the clique itself. We compute plausibility scores over test instances with correct predicted label, because local explanations extracted from wrong predictions are not reliable for analyzing model decisions. We report in Appendix E the corresponding accuracy scores of downstream classifiers.
Table 3 compares Plausibility scores where node features from different embeddings are used to train downstream predictors. Since typical embeddings exhibit semantic patterns distributed across many dimensions (Elhage et al., 2022), here we consider all the contributing dimensions employing a non-negative weighting function $f(*)=\mathrm{max(0,*)}$. This choice prevents bias toward methods that inherently produce disentangled semantics (e.g., by analyzing only top-ranked dimensions). DeepWalk, GraphAE, and InfWalk perform modestly, with DeepWalk scoring the highest among these on Ring-Cl and InfWalk showing relative strength on BA-Cl. Compared with earlier approaches, DGLFRM delivers a modest gain in Plausibility, most notably in link prediction, whereas GraphSAGE significantly underperforms across all tasks, especially in node classification. The addition of Dine improves both DeepWalk and GAE. DW+Dine excels with strong performance on Ring-Cl, SBM, and Tree datasets, while GAE+Dine achieves slightly worst results, particularly on node classification tasks, such as in Tr-Gr. Within the proposed models, DiSe-FCAE and DiSe-GAE consistently achieve the highest scores ranking as the best two methods overall.
In Table 4 we compare Plausibility for state-of-the-art local post-hoc explainers for graphs in node classification. We emphasize that our approach focuses on explaining model encodings, unlike methods such as GNNExplainer and PGExplainer, which explain model decisions. These methods present local explanation in the form of node and/or edge importance, whereas in our method, combined with feature-based explainer, the explanation format is a vector of feature importance, associated with a subgraph for each feature. To make a suitable comparison, we consider as the explanation presented by our method the subgraph associated to the most important embedding feature (according to the logistic classifier). Recalling Eq. (8), this approach is equivalent to choosing the function $f(\Psi_{j})=\mathbb{1}[j=\mathrm{argmax}\{\Psi_{1},\dots\Psi_{K}\}]$ to compute plausibility index. We observe GNNExplainer has uniform results across different input GNN models, instead PGExplainer performs best with GCN. DiSe-FCAE and DiSe-GAE outperforms the competitors in most of the cases, except with GCN+PGExplainer in ER-Cliques. Reported results show that our method is capable of producing subgraph-based local explanations with comparable, or even better, plausibility scores than GNNExplainer/PGExplainer. To enhance clarity, Appendix E includes qualitative examples that visualize the local explanations generated by the different methods.